Sanjana loves to read books(B) and likes to watch movies(M) as well. She derives double the satisfaction from reading a book than watching a movie. Her utility function is as follows

U(B, M) = 2B + M

Where B and M are the number of books and movies respectively.

a. Explain the shape of the indifference curve using diagrams. Word limit: 200]

b. The income available with her is INR60. And price of movie ticket is INR 10 and that of a book is INR 15. Will she spend her entire budget and why? What will be her consumption, explain why? [Word limit: 300



1
Expert's answer
2021-12-06T16:33:36-0500

a. Since books and movies are perfect isabstitutes, the the indifference curve is a linear downwarding sloping. Besides, from the utility function, Sanajana's preference for books is twice as her her preference for movies. therefore the slope of the indifference curve is is -2.


b. "Marginal rate of sabstitution (MRS) = \frac{MUB}{MUM} = 2"

"Slope of the budget line=\frac{15}{10} = 1.5"

"MRS> Slope of thebudget line"

Therefore she will only consume books

"Units of books consumed=\frac{60}{15}= 4 books"